OpenStudy (anonymous):

Passing through the point(5,-2) and parallel to x+5y+5=0. Write in general equation.

OpenStudy (kc_kennylau):

Find the slope first, slope = -A/B in Ax+By+C=0

OpenStudy (anonymous):

parallel => slope2=slope1 ffrom the irst equation: y=-1/5x-1 => slope 1 = -1/5. =>slope2=-1/5 => equation2: y=-1/5x+C, we don't know what is C, but we know that when x=5 y=-2 (passing through the point (5-2)), so -2=-1/5*5+C => C=-2+1=-1 => y=-1/5x-1 and in general equation: x+5y+5=0

OpenStudy (shamim):

ur required equation is y-y1=m(x-x1)

OpenStudy (shamim):

m=-1/5 x1=5 y1=-2

OpenStudy (anonymous):

true)

OpenStudy (anonymous):

I dont understand why y=-1/5x-1 equals to x+5y+5=0. Explain? :)

OpenStudy (shamim):

u know an equation of a line is y=mx+c here m= slope of a line

OpenStudy (shamim):

if u compare ur given line, u will get m=-1/5

OpenStudy (shamim):

@JC99

OpenStudy (anonymous):

@JC99, y=-1/5x-1 equals y+1/5x+1=0 yes? and it equals 5y+x+5=0 (multiplying by five) equals x+5y+5=0 that's why y=-1/5x-1 equals to x+5y+5=0 - these are just different records